MC ATC /AVC D=AR=MR $5.00 & $4.00 h PC $2.50 | 500

  1. Based on the level of output being produced, is this firm maximizing profit?  What is the dollar value of the profit being earned by the firm?  Use the lettering on the graph to identify the area of profit.
  2. How does the demand curve let you know this is a firm operating in perfect competition?
  3. What is the significance of Point E?  Point F?
  4. How much additional cost did the 500thunit add to total cost?  How do you know?
  5. Explain how the market will adjust to return this firm to long run equilibrium.  Be sure to discuss the changes that occur in the market and at the firm.  
Answer #1

1). Yes, it is the profit maximizing output, the perfectly competitive firm maximizes the profit where the marginal revenue equals the marginal cost.

Profit =\left ( P-ATC \right )\times Q

=\left ( 5-4\right )\times 500.

  500 The firm's economic profit equals $500.

The rectangle area 'Edch' equals the economic profit at that output.

2). In the perfect competition there is identical products , this makes the firms as the price takers. The does not have any market power , so the demand would be perfectly elastic and the perfectly elastic demand curve is a horizontal straight line parallel to the 'X' axis, and the demand is also equal to the marginal and the average revenue. In the perfect competition the demand should be perfectly elastic.

3). The point 'f' is the shutdown point ,and its is where the price equals the average variable cost. If the price goes below this point the firm should shutdown production and if the price is above this point the firm can continue production in the short run and not in the long run.

4). At point 'e' the firm earns zero economic profit , that is the total revenue will be equal to the total cost. It is the long run equilibrium of a perfectly competitive firm, the firms would be earning zero economic profit in the long run.

5). This is we call as the marginal cost, the marginal cost is the addition made to the total cost when additional unit of the commodity or service produced. At the 500th unit the profit is maximized, profits are maximized when the marginal cost is equal to the marginal revenue so the additional cost or the marginal cost at the 500 th unit would be $ 5.

6). At the 500th unit the firm is earning a positive economic profit, so other firms would be attracted by these positive profits so they will come into the market since there is no restriction on entry and exit. When the new firms come into the market that will decrease the demand demand for the existing firms so the demand decrease. The new firms will keep coming until the economic profits equals zero.